Frozen (quarter- pound) hamburger patties cost $5.80 for 12. How many can be purchased for $7.00?





Frozen (quarter- pound) hamburger patties cost $5.80 for 12. How many can be purchased for $7.00?

Answer by jayanthihemkos@yahoo.com(87)

Frozen (quarter- pound) hamburger patties cost $5.80 for 12. How many can be purchased for $7.00?
Cost of 12 hamburger patties is $ 5.80
Cost of 1 hamburger patties is 5.80/12 = $0.483
$0.483 gives 1 hamburger patties
$7 = ?
7/0.483
= 14.5 hamburger
i.e We can buy 14 and half hamburgers from $7

130% OF 50 = 130/100*50
= 65
New number now = 50+65 = 115